### v1

@@ Abstract
We revisit the efficacy of several practical methods for approximate machine unlearning developed for large-scale deep learning. In addition to complying with data deletion requests, one often-cited potential application for unlearning methods is to remove the effects of training on poisoned data. We experimentally demonstrate that, while existing unlearning methods have been demonstrated to be effective in a number of evaluation settings (e.g., alleviating membership inference attacks), they fail to remove the effects of data poisoning, across a variety of types of poisoning attacks (indiscriminate, targeted, and a newly-introduced Gaussian poisoning attack) and models (image classifiers and LLMs); even when granted a relatively large compute budget. In order to precisely characterize unlearning efficacy, we introduce new evaluation metrics for unlearning based on data poisoning. Our results suggest that a broader perspective, including a wider variety of evaluations, are required to avoid a false sense of confidence in machine unlearning procedures for deep learning without provable guarantees. Moreover, while unlearning methods show some signs of being useful to efficiently remove poisoned datapoints without having to retrain, our work suggests that these methods are not yet “ready for prime time,” and currently provide limited benefit over retraining.
@@ 1 Introduction
ModernMP and JZD have equal first-author contributions, and GK, AS, and SN have equal advisory contributions. Machine Learning (ML) models are often trained on large-scale datasets, which can include significant amounts of sensitive or personal data. This practice raises privacy concerns as the models can memorize and inadvertently reveal information about individual points in the training set. Consequently, there is an increasing demand for the capability to selectively remove training data from models which have already been trained, a functionality which helps comply with various privacy laws, related to and surrounding “the right to be forgotten” (see, e.g., the European Union’s General Data Protection Regulation (GDPR) (General Data Protection Regulation, ), the California Consumer Privacy Act (CCPA), and Canada’s proposed Consumer Privacy Protection Act (CPPA)). This functionality is known as machine unlearning (Cao and Yang, 2015), a field of research focused on "removing" specific training data points from a trained model upon request. The ideal goal is to produce a model that behaves as if the data was never included in the training process, effectively erasing all direct and indirect traces of the data. Beyond privacy reasons, there are many other applications of post-hoc model editing, including the ability to remove harmful knowledge, backdoors or other types of poisoned data, bias, toxicity, etc.
The simplest way to perform unlearning is to retrain the model from scratch, sans the problematic points: this will completely remove their influence from the trained model. However, this is often impractical, due to the large scale of modern ML systems. Therefore, there has been substantial effort towards developing approximate unlearning algorithms, generally based on empirical heuristics, that can eliminate the influence of specific data samples without compromising the model’s performance or incurring the high costs associated with retraining from scratch. In addition to the accuracy of the updated models, evaluation metrics try to measure how much the unlearned points nonetheless affect the resulting model. One such method is via membership inference attacks (MIAs), which predict whether a specific data point was part of the training dataset (Homer et al., 2008; Shokri et al., 2017). Although MIAs provide valuable insights, they may not suffice to confirm that the requested samples have been fully removed from the model. Since MIAs against deep learning models are themselves heuristics, and known MIAs can be computationally expensive to implement themselves Carlini et al. (2022a), even if a MIA suggests that a datapoint has been successfully unlearned, this does not guarantee that residual traces of the data do not remain, potentially allowing adversaries to recover sensitive information.
@@
Our high-level contributions are as follows:
@@ 1. $\bullet$
    Failure of current state-of-the-art unlearning algorithms: We evaluate seven state-of-the-art unlearning algorithms explored in machine unlearning literature, across standard language and vision classification tasks, in terms of their ability to mitigate the effects of data poisoning. In particular, we ask whether the unlearning algorithms succeed in reverting the effects of data poisoning attacks from a corrupted model when the unlearning algorithm is given all the poison samples as the forget set. Experimentally, we evaluate machine unlearning using indiscriminate, targeted, and Gaussian data poisoning attacks and show that (a) none of the current state-of-the-art unlearning algorithms can mitigate all of these data poisoning attacks, and (b) different data poisoning methods introduce different challenges for unlearning, and (c) the success of an unlearning method depends on the underlying task.
@@ 2. $\bullet$
    Introduction of a new evaluation measure: We introduce a new measure to evaluate machine unlearning based on Gaussian noise. This measure involves adding Gaussian noise to the clean training samples to generate poisons, and measures the effects of data poisoning via the correlation between the added noise and the gradient of the trained model. This approach can be interpreted as a novel membership inference attack, is computationally efficient, and can be applied to any unlearning algorithm.
@@ 3. $\bullet$
    Advocating for careful unlearning evaluation: By demonstrating that heuristic methods for unlearning can be misleading, we advocate for proper evaluations or provable guarantees for machine unlearning algorithms as the way forward.
@@ 2 Related Works
Machine unlearning. At this point, there exists a vast literature on machine unlearning (Cao and Yang, 2015), we focus on the most relevant subset here. Many works focus on removing the influence of training on a particular subset of points from a trained model (Ginart et al., 2019b; Wu et al., 2020; Golatkar et al., 2020a, b; Bourtoule et al., 2021; Izzo et al., 2021; Neel et al., 2021; Sekhari et al., 2021; Jang et al., 2022; Huang and Canonne, 2023; Wang et al., 2023). Others instead try to remove a subset of concepts (Ravfogel et al., 2022a, b; Belrose et al., 2023). In general, the goal is to excise said information without having to retrain the entire model from scratch. Some works focus on exactly unlearning (see, e.g., Bourtoule et al. (2021)), whereas others try to only approximately unlearn (e.g., Ginart et al. (2019a); Sekhari et al. (2021); Neel et al. (2021)), using a definition inspired by differential privacy (Dwork et al., 2006). Much of the work in this line focuses on unlearning in the context of image classifiers (e.g., Golatkar et al. (2020a); Goel et al. (2022); Kurmanji et al. (2023); Ravfogel et al. (2022a, b); Belrose et al. (2023)). Some approximate unlearning methods are general-purpose, using methods like gradient ascent (Neel et al., 2021), or are specialized for individual classes such as linear regression (Cook and Weisberg, 1980; Guo et al., 2019; Izzo et al., 2021) or kernel methods (Zhang and Zhang, 2021).
@@
Evaluating machine unlearning. Some of the works mentioned above focus on provable machine unlearning (either exact or approximate). That is, as long as the algorithm is carried out faithfully, the resulting model is guaranteed to have unlearned the pertinent points. However, many unlearning methods are heuristic, without provable guarantees. Alternatively, we may be given access to an unlearning procedure as a black box. In either case, we may want to measure or audit how well an unlearning method performed. Several works (see, e.g., Kurmanji et al. (2024); Goel et al. (2022); Golatkar et al. (2020a, b); Graves et al. (2021); Ma et al. (2022); Pawelczyk et al. (2023, 2024); Hayes et al. (2024)) mostly perform various adaptations of membership inference attacks to the unlearning setting. However, essentially all of these methods search for “direct” influence of a training point on the resulting model: that is, how the trained model responds to the particular point that was unlearned. In contrast, our work complements such techniques, by measuring removal of indirect influence of a point on the resulting model, via data poisoning attacks. Our results show that even if machine unlearning methods appear effective at removing direct influence, they may not be effective at removing indirect influence.
@@
Sommer et al. (2022) proposed a verification framework for machine unlearning by adding backdoor triggers to the training dataset, however, they do not perform any evaluations for the current state-of-the-art machine unlearning algorithms proposed for deep learning scenarios. Goel et al. (2024) asks whether machine unlearning can mitigate the effects of data poisoning when the unlearning algorithm is only given an incomplete subset of the poison samples. On the other hand, we employ stronger attacks which result in showing that machine unlearning is in fact unable to remove the influence of data poisoning. Our work thus complements these prior works by designing novel clean-label data poisoning methods such as Gaussian data poisoning, and extensive evaluation on practically used state-of-the-art machine unlearning algorithms
Data poisoning attacks. In a data poisoning attack, an adversary may introduce or modify a small portion of the training data, and their goal is to elicit some undesirable behavior in a model trained on said data. One type of attack is a targeted data poisoning attack (Koh and Liang, 2017; Shafahi et al., 2018; Huang et al., 2020; Guo and Liu, 2020; Aghakhani et al., 2021), in which the goal is to cause a model to misclassify a specific point in the test set. Another type of attack is an untargeted (or indiscriminate) data poisoning attack (Biggio et al., 2012; Muñoz-González et al., 2017; Steinhardt et al., 2017; Koh et al., 2022; Lu et al., 2022, 2023), wherein the attacker seeks to reduce the test accuracy as much as possible. Though we do not focus on them in our work, there also exist backdoor attacks (Gu et al., 2017), in which training points are poisoned with a backdoor pattern, such that test points including the same pattern are misclassified.
Poisoning machine unlearning systems. An orthogonal line of work investigates data poisoning attacks against machine unlearning pipelines (see, e.g., Chen et al. (2021); Marchant et al. (2022); Carlini et al. (2022b); Di et al. (2023); Qian et al. (2023); Liu et al. (2024)). These works generally show that certain threats can arise even if unlearning is performed with provable guarantees, whereas we focus on data poisoning threats in standard (i.e., not machine unlearning) pipelines, that ought to be removed by an effective machine unlearning procedure (in particular, they would be removed by retraining from scratch without the poisoned points).
3 Machine Unlearning: Preliminaries and Algorithms
@@
We formalize the machine unlearning setting and introduce relevant notation. Let $S_{\textsf{train}}$ and $S_{\textsf{test}}$ be training and test datasets for an ML model, respectively, each consisting of samples of the form $(x,y)$ where $x\in\mathbb{R}^{d}$ denotes the covariate (e.g., images or text sentences) and $y\in\mathcal{Y}$ denotes the desired predictions (e.g., labels or text predictions). The unlearner starts with a model $\theta_{\textsf{initial}}$ obtained by running a learning algorithm on the training dataset $S_{\textsf{train}}$; the model $\theta_{\textsf{initial}}$ is trained to have small loss over the training dataset, and by proxy, the test dataset as well. Given a set of deletion requests $U\subseteq S_{\textsf{train}}$, the unlearner runs an unlearning algorithm to update the initial trained model $\theta_{\textsf{initial}}$ to an updated model $\theta_{\textsf{updated}}$, with the goal that (a) $\theta_{\textsf{updated}}$ continues to perform well on the test dataset $S_{\textsf{test}}$, and (b) $\theta_{\textsf{updated}}$ does not have any influence of the delete set $U$.
The simplest method for eliminating the samples $U$ from $\theta_{\textsf{initial}}$ is "retraining from scratch": delete $U$ from $S_{\textsf{train}}$ and then run the learning algorithm again on the remaining data $S_{\textsf{train}}\setminus U$. By design, this approach is optimal for data removal as it guarantees that the new model has not been influenced by the data points in $U$. Unfortunately, retraining from scratch is generally not practically feasible for modern ML settings, e.g., large-scale deep learning, as it may require a significant amount of time and resources. Consequently, much of the research in machine unlearning has been directed towards developing approximate unlearning methods, often without rigorous theoretical guarantees, that can update $\theta_{\textsf{initial}}$ in a computationally- and resource-efficient manner to remove the effects of $U$ [^1] While we present use these algorithms in the batch unlearning setting, consisting a single stage of learning and unlearning only, we remark that all of these algorithms can be extended to the iterative / online unlearning setting. We list some of the most popular approximate unlearning methods below with details deferred to Appendix A.1. .
  1. $\bullet$
    Gradient Descent (GD) (Neel et al., 2021): GD continues to train the model $\theta_{\textsf{initial}}$ on the remaining dataset $S_{\textsf{train}}\setminus U$ by using gradient descent. In particular, we obtain $\theta_{\textsf{updated}}$ via
    $\theta_{t+1}\leftarrow\theta_{t}-\eta g_{t}(\theta_{t})\qquad\text{with}\qquad\theta_{1}=\theta_{\textsf{initial}},$
    where $\eta$ denotes the step size and $g_{t}$ denotes a (mini-batch) gradient computed for the the training loss $\widehat{\operatorname{\mathbb{E}}}_{(x,y)\in S_{\textsf{train}}\setminus U}\left[\ell((x,y),\theta)\right]$ defined using the remaining dataset $S_{\textsf{train}}\setminus U$, where $\ell$ is a loss function, e.g.,  cross-entropy loss, hinge loss, etc.
  2. $\bullet$
    Noisy Gradient Descent (NGD) (Chien et al., 2024; Chourasia and Shah, 2023): NGD is a simple modification of GD where we obtain $\theta_{\textsf{updated}}$ via the update
    $\theta_{t+1}\leftarrow\theta_{t}-\eta(g_{t}(\theta_{t})+\xi_{t})\qquad\text{with}\qquad\theta_{1}=\theta_{\textsf{initial}},$
    where $\xi_{t}\sim\mathcal{N}(0,\sigma^{2})$ denotes an independently sampled Gaussian noise and $g_{t}$ denotes a (mini-batch) gradient computed for the training loss defined using the remaining dataset $S_{\textsf{train}}\setminus U$.
  3. $\bullet$
    Gradient Ascent (GA) (Graves et al., 2021; Jang et al., 2022): GA is an unlearning algorithm which attempts to remove the influence of the forget set $U$ from the trained model by simply reversing the gradient updates that contain information about $U$. In particular, we update via
    $\theta_{t+1}\leftarrow\theta_{t}+\eta g_{t}(\theta_{t})\qquad\text{with}\qquad\theta_{1}=\theta_{\textsf{initial}},$
    where $g_{t}$ denotes a (mini-batch) gradient computed on the loss $\widehat{\operatorname{\mathbb{E}}}_{(x,y)\in U}\left[\ell((x,y),\theta)\right]$ on the deletion set.
  4. $\bullet$
    EUk (Goel et al., 2022): Exact Unlearning the last k layers (EUk) is an unlearning approach for deep learning settings that simply retrains from scratch the last k layers (that are closest to the output/prediction layer) of the neural network, while keeping all previous layers fixed.
  5. $\bullet$
    CFk (Goel et al., 2022): Catastrophically forgetting the last k layers (CFk) is a straightforward modification of EUk, with the only difference being that instead of retraining from scratch, we continue training the weights in the last k layers on the retain set $S_{\textsf{train}}\setminus U$.
  6. $\bullet$
    SCRUB (Kurmanji et al., 2024): SCalable Remembering and Unlearning unBound (SCRUB) is a state-of-the-art unlearning method for deep learning settings. It casts the unlearning problem into a student-teacher framework, and computes the parameter $\theta_{\textsf{updated}}$ by minimizing the objective
    ${\small\widehat{\operatorname{\mathbb{E}}}_{(x,y)\sim S_{\textsf{train}}\setminus U}\left[\mathrm{KL}{\left(M_{\theta_{\textsf{initial}}}(x)\|M_{\theta}(x)\right)}+\ell(\theta;(x,y))\right]-\widehat{\operatorname{\mathbb{E}}}_{(x,y)\sim U}\left[\mathrm{KL}{\left(M_{\theta_{\textsf{initial}}}(x)\|M_{\theta}(x)\right)}\right]}$
  7. $\bullet$
    NegGrad+ (Kurmanji et al., 2024): NegGrad+ is a finetuning based unlearning approach, and computes $\theta_{\textsf{updated}}$ by minimizing the objective
    using gradient descent, where $\beta\in(0,1)$ is a hyperparameter.
While all of the above algorithms are designed to retain performance on the remaining training dataset $S_{\textsf{train}}\setminus U$ and the test dataset $S_{\textsf{test}}$, prior works also evaluated the unlearning capability of these methods using various heuristics. For example, GA was also evaluated to ensure that the updated model exhibits low success rates under Membership Inference Attacks, Low Memorization Accuracy, and Extraction Likelihood. Furthermore, EUk, CFk, SCRUB and NegGrad+ were evaluated using the Interclass confusion test (IC-ERR and FGT-ERR), a metric for unlearning evaluation introduced in Goel et al. (2022). Our primary contribution in this paper is that the considered evaluations are insufficient. In the following section, we show via experiments using data poisoning methods that the above-listed state-of-the-art machine unlearning algorithms do not succeed in fully removing all the influence of the deletion set $U$ from the updated model $\theta_{\textsf{updated}}$.

### v2

@@ Abstract
We revisit the efficacy of several practical methods for approximate machine unlearning developed for large-scale deep learning. In addition to complying with data deletion requests, one often-cited potential application for unlearning methods is to remove the effects of poisoned data. We experimentally demonstrate that, while existing unlearning methods have been demonstrated to be effective in a number of settings, they fail to remove the effects of data poisoning across a variety of types of poisoning attacks (indiscriminate, targeted, and a newly-introduced Gaussian poisoning attack) and models (image classifiers and LLMs); even when granted a relatively large compute budget. In order to precisely characterize unlearning efficacy, we introduce new evaluation metrics for unlearning based on data poisoning. Our results suggest that a broader perspective, including a wider variety of evaluations, are required to avoid a false sense of confidence in machine unlearning procedures for deep learning without provable guarantees. Moreover, while unlearning methods show some signs of being useful to efficiently remove poisoned data without having to retrain, our work suggests that these methods are not yet “ready for prime time,” and currently provide limited benefit over retraining.
@@ 1 Introduction
Modern Machine Learning (ML) models are often trained on large-scale datasets, which can include significant amounts of sensitive or personal data. This practice raises privacy concerns as the models can memorize and inadvertently reveal information about individual points in the training set. Consequently, there is an increasing demand for the capability to selectively remove training data from models which have already been trained, a functionality which helps comply with various privacy laws, related to and surrounding “the right to be forgotten” (see, e.g., the European Union’s General Data Protection Regulation (GDPR, ), the California Consumer Privacy Act (CCPA), and Canada’s proposed Consumer Privacy Protection Act (CPPA)). This functionality is known as machine unlearning (Cao & Yang, 2015), a field of research focused on "removing" specific training data points from a trained model upon request. The goal is to produce a model that behaves as if the data was never included in the training process, effectively erasing all direct and indirect traces of the data. Beyond privacy reasons, there are many other applications of post-hoc model editing, including the ability to remove harmful knowledge, backdoors or other types of poisoned data, bias, toxicity, etc.
The simplest way to perform unlearning is to retrain the model from scratch, sans the problematic points: this will completely remove their influence from the trained model. However, this is often impractical, due to the large scale of modern ML systems. Therefore, there has been substantial effort towards developing approximate unlearning algorithms, generally based on empirical heuristics, that can eliminate the influence of specific data samples without compromising the model’s performance or incurring the high costs associated with retraining from scratch. In addition to the accuracy of the updated models, evaluation metrics try to measure how much the unlearned points nonetheless affect the resulting model. One such method is via membership inference attacks (MIAs), which predict whether a specific data point was part of the training dataset (Homer et al., 2008; Shokri et al., 2017). Although MIAs provide valuable insights existing unlearning MIAs are computationally expensive to implement themselves (Pawelczyk et al., 2024; Hayes et al., 2024; Kurmanji et al., 2024). Even if a MIA suggests that a datapoint has been successfully unlearned, this does not guarantee that residual traces of the data do not remain, potentially allowing adversaries to recover sensitive information.
@@
In this work, we evaluate eight state-of-the-art unlearning algorithms explored in machine unlearning literature, across standard language and vision tasks, in terms of their ability to mitigate the effects of data poisoning. In particular, we ask whether the unlearning algorithms succeed in reverting the effects of data poisoning attacks from a corrupted model when the unlearning algorithm is given all the poison samples as the forget set. Our high-level contributions are as follows:
@@ 1. $\bullet$
    Failure of current state-of-the-art unlearning algorithms: We stress test machine unlearning using indiscriminate, targeted, backdoor, and Gaussian data poisoning attacks and show that i) none of the current state-of-the-art unlearning algorithms can mitigate all of these data poisoning attacks, ii) different data poisoning methods introduce different challenges for unlearning, and iii) the success of an unlearning method depends on the underlying task.
@@ 2. $\bullet$
    Introduction of a new evaluation measure: We introduce a new measure to evaluate machine unlearning based on Gaussian noise. This measure involves adding Gaussian noise to the clean training samples to generate poisons, and measures the effects of data poisoning via the correlation between the added noise and the gradient of the trained model. This approach can be interpreted as a novel membership inference attack, is computationally efficient, compatible across all data domains (tabular, image, language) and can be applied to any unlearning algorithm. We release the code for our Gaussian data poisoning method at: https://github.com/MartinPawel/OpenUnlearn.
@@ 3. $\bullet$
    Insights into Unlearning Failure: We develop and experimentally validate two novel hypotheses explaining why unlearning methods fail under data poisoning attacks.
  4. $\bullet$
    Advocating for detailed unlearning evaluation: By demonstrating that heuristic methods for unlearning can be misleading, we advocate for proper evaluations or provable guarantees for machine unlearning algorithms as the way forward.
@@ 2 Related Works
Machine unlearning. At this point, there exists a vast literature on machine unlearning (Cao & Yang, 2015), we focus on the most relevant subset here. Many works focus on removing the influence of training on a particular subset of points from a trained model (Ginart et al., 2019; Wu et al., 2020; Golatkar et al., 2020a; b; Bourtoule et al., 2021; Izzo et al., 2021; Neel et al., 2021; Sekhari et al., 2021; Jang et al., 2022; Wang et al., 2023). Others instead try to remove a subset of concepts (Ravfogel et al., 2022a; b; Belrose et al., 2023). In general, the goal is to excise said information without having to retrain the entire model from scratch. Some works focus on exactly unlearning (see, e.g., Bourtoule et al. (2021)), whereas others try to only approximately unlearn (e.g., Sekhari et al. (2021); Neel et al. (2021)). Much of the work in this line focuses on unlearning in the context of image classifiers (e.g., Golatkar et al. (2020a); Goel et al. (2022); Kurmanji et al. (2023); Ravfogel et al. (2022a; b); Belrose et al. (2023); Fan et al. (2023); Chen et al. (2024)). Some approximate unlearning methods are general-purpose, using methods like gradient ascent (Neel et al., 2021), or are specialized for individual classes such as linear regression (Guo et al., 2019; Izzo et al., 2021) or kernel methods (Zhang & Zhang, 2021).
@@
Evaluating machine unlearning. Some of the works mentioned above focus on provable machine unlearning (either exact or approximate). That is, as long as the algorithm is carried out faithfully, the resulting model is guaranteed to have unlearned the pertinent points. However, many unlearning methods are heuristic, without provable guarantees. This is why we may want to measure or audit how well an unlearning method performed. Several works (see, e.g., Kurmanji et al. (2024); Goel et al. (2022); Golatkar et al. (2020a; b); Graves et al. (2021); Ma et al. (2022); Pawelczyk et al. (2023; 2024); Hayes et al. (2024)) mostly perform various adaptations of membership inference attacks to the unlearning setting that either suffer from low statistical power (Kurmanji et al., 2024; Golatkar et al., 2020b; Graves et al., 2021) or require training hundreds of shadow models to evaluate unlearning (Pawelczyk et al., 2024; Kurmanji et al., 2023; Hayes et al., 2024). Relative to these works, our Gaussian poisoning attack has high statistical power at low false positive rates and can be cheaply run using one training run of the model. Sommer et al. (2022) proposed a verification framework for machine unlearning by adding backdoor triggers to the training dataset, however, they do not perform any evaluations for the current state-of-the-art machine unlearning algorithms. Goel et al. (2024) asks whether machine unlearning can mitigate the effects of data poisoning when the unlearning algorithm is only given an incomplete subset of the poison samples.
@@
Compared to these works, we employ stronger attacks which result in showing that machine unlearning is in fact unable to remove the influence of data poisoning. Our work thus complements these prior works by designing novel clean-label data poisoning methods such as Gaussian data poisoning, and extensive evaluation on practically used state-of-the-art machine unlearning algorithms
3 Machine Unlearning Evaluation Preliminaries
@@
We formalize the machine unlearning setting and introduce relevant notation. Let $S_{\textsf{train}}$ and $S_{\textsf{test}}$ be training and test datasets for an ML model, respectively, each consisting of samples of the form $z=(x,y)$ where $x\in\mathbb{R}^{d}$ denotes the covariate (e.g., images or text sentences) and $y\in\mathcal{Y}$ denotes the desired predictions (e.g., labels or text predictions). The unlearner starts with a model $\theta_{\textsf{initial}}$ obtained by running a learning algorithm on the training dataset $S_{\textsf{train}}$; the model $\theta_{\textsf{initial}}$ is trained to have small loss over the training dataset, and by proxy, the test dataset as well. Given a set of deletion requests $U\subseteq S_{\textsf{train}}$, the unlearner runs an unlearning algorithm to update the initial trained model $\theta_{\textsf{initial}}$ to an updated model $\theta_{\textsf{updated}}$, with the goal that i) $\theta_{\textsf{updated}}$ continues to perform well on the test dataset $S_{\textsf{test}}$, and ii) $\theta_{\textsf{updated}}$ does not have any influence of the delete set $U$. Our focus in this paper is to evaluate whether a given unlearning algorithm satisfies these desiderata.
[…]
Prior approach for machine unlearning evaluation via membership inference. Machine unlearning is typically evaluated by checking if an instance $z$ is a member of the training set (MEMBER) by checking if the loss $\ell$ of the model $\theta_{\textsf{updated}}$ is lower than or equal to a threshold $\tau_{L}$ (Shokri et al., 2017):
$M_{\text{Loss}}(z)=\begin{cases}\emph{{MEMBER}}&\text{ if }\ell(\theta_{\textsf{updated}},z)\leq\tau_{L}\\ \emph{{NON-MEMBER}}&\text{ else}.\end{cases}$ (1)
Under exact unlearning, this attack should have trivial accuracy, achieving a true positive rate that equals the false positive rate (i.e., TPR = FPR) at every value of $\tau_{L}$. The unlearning error is then measured by the extent to which the classifier achieves nontrivial accuracy when deciding whether samples are MEMBER or NON-MEMBER, in particular focusing on the tradeoff curve between TPR at FPR at or below 0.01 denoted as TPR@@@@FPR=0.01 (Carlini et al., 2022a).